# Deep Discrete Encoders: Identifiable Deep Generative Models for Rich Data with Discrete Latent Layers

### v1

@@ 7 Discussion
where Unif$(a,b)$ denotes the uniform distribution on the interval $(a,b)$.
ParFam Accuracy($\mathcal{G}$) RMSE$(\boldsymbol{\Theta})$ Time (s) # iterations
Initialization \$N$ 1000 4000 1000 4000 1000 4000 1000 4000
@@
Table S.1: Accuracy measures for $\mathcal{G}$ and $\boldsymbol{\Theta}$, computation time and iterations for 2-layer DDE estimates under different initializations. For the Accuracy$(\mathcal{G})$ column, larger is better. For the other columns, smaller is better.
@@
S.3.1 Detailed Update Formulas
Simplified M-step for the PEM Algorithm.
We display more explicit expressions for the M-step in Algorithm 1. Recall that we need to solve three optimizations detailed in equations (8)-(10). By plugging-in the distributional formulas from Section 2, we can simplify each equation as follows. The computations are straightforward and we just display the simplified formulas.
First, the maximization in (8) simplifies to updating
$p_{l}^{[t+1]}=\frac{\sum_{i,\operatorname{\boldsymbol{\alpha}^{(1)}},\operatorname{\boldsymbol{\alpha}^{(2)}}2}\alpha_{l}^{(2)}\varphi^{[t+1]}_{i,\operatorname{\boldsymbol{\alpha}^{(1)}},\operatorname{\boldsymbol{\alpha}^{(2)}}2}}{N},$
for $l\in[K^{(2)}]$. The maximization in (9) simplifies to maximizing the $k$th row of $\mathbf{B}^{(2)}$ for each $k\in[K^{(1)}]$:
${\boldsymbol{\beta}}_{k}^{(2),[t+1]}$
$=$ $\underset{{\boldsymbol{\beta}}^{(2)}\in\mathbb{R}^{K^{(2)}+1}}{\operatorname*{argmax}}\sum_{i,\operatorname{\boldsymbol{\alpha}^{(1)}},\operatorname{\boldsymbol{\alpha}^{(2)}}2}\log\mathbb{P}(\mathbf{A}_{i,k}=\alpha_{k}^{(1)};g_{\text{logistic}}(\beta_{0}^{(2)}+\sum_{l=1}^{K^{(2)}}\alpha_{l}^{(2)}\beta_{l}^{(2)}))\varphi_{i,\operatorname{\boldsymbol{\alpha}^{(1)}},\operatorname{\boldsymbol{\alpha}^{(2)}}2}^{[t+1]}-p_{\lambda_{2},\tau}({\boldsymbol{\beta}}^{(2)}).$
Finally, the maximization in (10) boils down to maximizing the $j$th row of $\mathbf{B}^{(1)}$ and $\gamma_{j}$ for each $j\in[J]$:
$({\boldsymbol{\beta}}_{j}^{(1),[t+1]},\gamma_{j}^{[t+1]})$
$=$ $\underset{{\boldsymbol{\beta}}^{(1)}\in\mathbb{R}^{K^{(1)}+1},\gamma>0}{\operatorname*{argmax}}~{}\sum_{i,\operatorname{\boldsymbol{\alpha}^{(1)}},\operatorname{\boldsymbol{\alpha}^{(2)}}2}\log\mathbb{P}(Y_{i,j};\beta_{0}^{(1)}+\sum_{k=1}^{K^{(1)}}\alpha_{k}^{(1)}\beta_{k}^{(1)},\gamma)\varphi_{i,\operatorname{\boldsymbol{\alpha}^{(1)}},\operatorname{\boldsymbol{\alpha}^{(2)}}2}^{[t+1]}-p_{\lambda_{1,N},\tau}({\boldsymbol{\beta}}^{(1)}),$
@@
S.3.2 Estimating the Number of Latent Variables
@@
    We postpone the details of the formula of EBIC$(k)$ to Section S.3.4. Here, we consider EBIC instead of more traditional information criteria such as AIC and BIC since it is designed to handle large parameter spaces.
@@
S.3.3 Implementation Details
We elaborate on the choices made to implement the EM Algorithms 1 and 3.
@@
Based on the consistency result in Theorem 3, we consider $\lambda_{N}=N^{1/4},\tau_{N}=3N^{-0.3}$ for the simulation studies. For real data analysis, tuning parameters were selected from the following grid:
@@
In terms of implementing the PEM algorithm to select the number of latent variables, $K^{(1)}$ or $K^{(2)}$, we consider a more generous threshold of $5003\times N/$ 500. This is for the sake of faster implementation, as we only use the resulting likelihood to implement the EBIC and LRT method.
@@
For practical implementation for a larger dataset, we recommend the user to modify the M-step to a faster but approximate optimization procedure. For example, one may choose to do a one-step gradient ascent in each iteration of the M-step.
Additional latent variable permutation for simulations.
@@
S.3.4 Definition of EBIC
Consider a 2-latent layer DDE with $\mathcal{K}=(K^{(1)},K^{(2)})$ latent variables. This is a parametric model with $|\boldsymbol{\Theta}|=J(K^{(1)}+1)+K^{(1)}(K^{(2)}+1)+K^{(2)}+J$ (if there exists a dispersion parameter) or $J(K^{(1)}+1)+K^{(1)}(K^{(2)}+1)+K^{(2)}$ (otherwise) parameters. Here, we do not count the discrete parameters $\mathcal{G}$, as they are implied by the coefficients $\mathcal{B}$. Let $\widehat{\boldsymbol{\Theta}}_{\mathcal{K}}$ and $\ell_{\mathcal{K}}(\widehat{\boldsymbol{\Theta}}_{\mathcal{K}};\mathbf{Y})$ respectively denote the MLE and the log-likelihood under the 2-layer DDE with $\mathcal{K}$ parameters. Also, let df$(\mathcal{K})$ be the number of non-zero parameters in $\widehat{\boldsymbol{\Theta}}_{\mathcal{K}}$. Then, following Chen and Chen, (2008), the EBIC objective in Section S.3.2 is defined as follows:
@@
S.4.1 Simulation Results Under the Generic Identifiable Parameters $\mathcal{B}_{g}$ and Computation Time
@@
We present the analogs of Figure 3 under the generically identifiable true parameters $\mathcal{B}_{g}$ defined as:
$\mathbf{B}_{g}^{(d)}=\begin{pmatrix}-2\mathbf{1}_{K^{(d)}}&\mathbf{B}_{2}^{(d)}\\ -4\mathbf{1}_{K^{(d)}}&\mathbf{B}_{2}^{(d)\top}\\ -2\mathbf{1}_{K^{(d)}}&\mathbf{B}_{1}^{(d)}\end{pmatrix},\quad\text{where}\quad\beta_{2;j,k}^{(d)}$ :=$:=\begin{cases}4&\text{if }k=j,\\ 4/3&\text{ else if }0<k-j\leq\lceil K^{(d)}/3\rceil,\\ 0&\text{otherwise},\end{cases}$ (S.31)
and $\mathbf{B}_{1}^{(d)}$ is the matrix defined in Equation 13. Note that $\mathbf{B}_{g}^{(d)}$ is a less sparse version of $\mathbf{B}_{s}^{(d)}$, where the identity matrices are modified to $\mathbf{B}_{2}^{(d)}$. While $\mathbf{B}_{2}^{(d)}$ and $\mathcal{B}_{g}$ have the same value of 4 on the main diagonals, $\mathcal{B}_{s}$ is sparser. In particular, $\mathcal{B}_{s}$ has two exclusive children per latent variable whereas $\mathcal{B}_{g}$ has none.
While all other simulation settings are identical to that described in the main text, we make the following two changes. First, we do not consider the largest parameter dimension of $(J,K^{(1)},K^{(2)})=(90,30,10)$ in this generically identifiable setting. This is because the layerwise initialization turned out to be less effective in this high-dimensional but less-sparse scenario. Second, for Poisson-based-DDEs, we modify the intercept values to be smaller compared to the values in (13) as follows. When $K^{(2)}=2$, we consider a smaller intercept parameter for $\mathbf{B}_{g}^{(1)}$ as follows:
@@
The results, presented in Tables S.3-S.5 (error of estimating $\mathcal{G}$) and S.6-S.8 (error of estimating $\boldsymbol{\Theta}$) in Section S.4.4, show that the PEM estimates exhibit a faster rate of convergence as $N$ increases. In particular, the RMSE values decrease at the parametric rate ${1}/{\sqrt{N}}$, as established in Theorem 3. We believe that the lower estimation accuracy under SAEM is due to multiple approximations within the SAEM algorithm such as approximate sampling and the objective function updates. We recommend using PEM instead of SAEM to estimate DDEs when the latent dimension is small.
@@
S.4.2 Experiments With Varying Numbers of Gibbs Samples in Algorithm 3
@@
ParFam Accuracy($\mathcal{G}$) RMSE$(\boldsymbol{\Theta})$ time (s) # iterations
# Gibbs $C$$N$ 1000 4000 1000 4000 1000 4000 1000 4000
@@
S.4.3 Simulations for Selecting the Number of Latent Variables
@@
Figure S.3: Selection accuracy of $K^{(1)}$ under the two-latent-layer DDE with true parameters $\mathcal{B}_{s}$. The spectral ratio estimator shows near-perfect accuracy for large $N$.
@@
We fit the three estimators $400$ times for each scenario, and display the correct selection percentage for the Bernoulli, Poisson, and Normal-based DDEs with true parameters $\mathcal{B}_{s}$/$\mathcal{B}_{g}$ in Figure S.3/Figure S.4. Comparing the three estimators, the spectral ratio-based estimator has near-perfect accuracy when $N$ is large enough, say $4000$. This demonstrates the empirical consistency of the spectral ratio estimator. In contrast, the consistency of the EBIC and LRT estimators is not clear for Bernoulli and Poisson responses [^2] To be more precise, as we are using approximate level $\alpha=0.01$ tests, the correct selection percentage of the LRT estimator should converge to 0.99.. Thus, we conclude that for a large enough $N$, it is desirable to select $K^{(1)}$ using the spectral ratio estimator. Among the three response types, the Bernoulli case with its nonlinear link function and limited response values is the most challenging. In contrast, the Normal case with a linear observed layer and continuous responses achieves near-perfect selection accuracy, which is consistent with earlier observations regarding parameter estimation accuracy.
@@
We also apply these three estimators to select $K^{(2)}$, assuming that $K^{(1)}=6$ is correctly estimated or known. The candidate set for $K^{(2)}$ is $\mathfrak{K}:=\{1,2,3\}$. Figure S.5/Figure S.6 displays the correct selection percentage under the true parameters $\mathcal{B}_{s}$/$\mathcal{B}_{g}$. Here, we implement the LRT estimator by naively assuming that Wilk’s theorem holds. The overall trends are similar to those observed in the previous figures for selecting $K^{(1)}$. Under the sparse true parameter $\mathcal{B}_{s}$, the spectral ratio estimator performs well across all response types and the LRT estimator has the lowest, but still decent accuracy. Interestingly, the EBIC estimator outperforms the spectral ratio estimator in more challenging scenarios, such as Bernoulli responses with small $N$ or under the Poisson responses with less sparse true parameters $\mathcal{B}_{g}$. Unlike the case for selecting $K^{(1)}$, it is not clear that the spectral ratio estimator outperforms the EBIC estimator for selecting $K^{(2)}$.
@@
S.4.4 Exact Values of Estimation Errors
@@
Tables S.3-S.5 display the average entrywise estimation accuracy for $\mathcal{G}$ under each parametric family. These values correspond to the upper rows of Figures 3 and S.1.
@@
Tables S.6S.8 reports the RMSE values for estimating continuous parameters $\boldsymbol{\Theta}$. These values correspond to the bottom rows of Figures 3 and S.1.
@@
Tables S.9-S.10 reports the accuracy of selecting $\mathcal{K}$ under Bernoulli and Poisson-based DDEs, which correspond to Figure S.3 and Figure S.4. Accuracy values for Normal-based DDEs are omitted, as all methods demonstrated near-perfect accuracy in this setting. Similarly, Tables S.11-S.12 display the results for selecting $K^{(2)}$, which corresponds to Figures S.5 and S.6.
@@
ParFam, value $N$ Method \ $\widehat{K}^{(1)}$ 4 5 6 7 8
Bernoulli, $\mathcal{B}_{s}$ 500 EBIC 0.25 96.25 3.5
LRT 0.5 98.75 0.75
Spectral 1.75 30.0 67.5 0.75
1000 EBIC 94.5 5.5
LRT 96.5 3.5
Spectral 0.25 18.25 81.5
2000 EBIC 88.25 11.75
LRT 93.5 6.25 0.25
Spectral 13.0 87.0
4000 EBIC 86.0 12.25 1.75
LRT 90.25 9.25 0.5
Spectral 3.75 96.25
6000 EBIC 86.0 13.5 1.5
LRT 87.75 12.25
Spectral 0.5 99.5
8000 EBIC 87.75 11.0 1.25
LRT 88.25 10.75 1.00
Spectral 0.75 99.25
Bernoulli, $\mathcal{B}_{g}$ 500 EBIC 84.5 13.75 1.75
LRT 0.75 1.0 86.25 12.0
Spectral 8.5 4.25 87.25
1000 EBIC 87.25 10.5 2.5
LRT 80.5 18.5 1.0
Spectral 0.5 99.5
2000 EBIC 85.25 11.5 3.25
LRT 82.75 16.25 1.0
Spectral 100.0
4000 EBIC 97.0 3.0
LRT 92.25 7.75
Spectral 100.0
6000 EBIC 97.5 2.5
LRT 91.5 8.5
Spectral 100.0
8000 EBIC 98.0 2.0
LRT 90.75 9.25
Spectral 100.0
@@
ParFam, value $N$ Method \ $\widehat{K}^{(1)}$ 4 5 6 7 8
Poisson, $\mathcal{B}_{s}$ 500 EBIC 99.5 0.5
LRT 99.5 0.5
Spectral 100.0
1000 EBIC 99.0 1.0
LRT 99.0 1.0
Spectral 100.0
2000 EBIC 97.5 2.5
LRT 96.25 3.75
Spectral 100.0
4000 EBIC 98.25 1.5 0.25
LRT 97.25 2.75
Spectral 100.0
6000 EBIC 99.0 1.0
LRT 97.25 2.75
Spectral 100.0
8000 EBIC 98.75 0.75 0.5
LRT 97.75 1.75 0.5
Spectral 100.0
Poisson, $\mathcal{B}_{g}$ 500 EBIC 92.0 7.75 0.25
LRT 1.0 91.0 8.0
Spectral 8.5 91.5
1000 EBIC 90.0 9.5 0.5
LRT 88.5 10.75 0.75
Spectral 2.75 97.25
2000 EBIC 90.5 8.5 1.0
LRT 0.25 87.25 11.5 1.0
Spectral 0.5 99.5
4000 EBIC 91.25 7.25 1.5
LRT 89.0 9.25 1.75
Spectral 100.0
6000 EBIC 95.0 4.5 0.5
LRT 91.25 7.75 1.0
Spectral 100.0
8000 EBIC 94.0 5.0 1.0
LRT 90.75 8.25 1.0
Spectral 100.0
@@
ParFam, value $N$ Method \ $\widehat{K}^{(2)}$ 1 2 3
Bernoulli, $\mathcal{B}_{s}$ 500 EBIC 0.5 99.5

### v2

@@ 7 Discussion
where Unif(a,b)(a,b) denotes the uniform distribution on the interval $(a,b)$.
[…]
@@
[…]
@@
S.3.1 Penalized EM Algorithm
This supplement presents the details of the standard penalized EM algorithm (PEM) that was mentioned in the main text, which does not use stochastic approximation. The PEM computes the penalized maximum likelihood estimator in (6) to estimate the potentially sparse coefficients $\mathbf{B}^{(1)}$ and $\mathbf{B}^{(2)}$. The PEM is an iterative procedure that consists of an expectation step followed by a penalized maximization step (Green, , 1990; Chen et al., , 2015). In the $(t+1)$th iteration, the E-step computes the expectation of the complete data penalized-log-likelihood
${\boldsymbol{p}}^{[t+1]}:=$ $\operatorname*{argmax}_{{\boldsymbol{p}}}\sum_{i=1}^{N}\mathbb{E}\left[\log\mathbb{P}(\mathbf{A}_{i}^{(2)};{\boldsymbol{p}});{\boldsymbol{p}}^{[t]}\right],$ (S.32)
$\mathbf{B}^{(2),[t+1]}:=$ $\operatorname*{argmax}_{\mathbf{B}^{(2)}}\sum_{i=1}^{N}\mathbb{E}\left[\log\mathbb{P}(\mathbf{A}_{i}^{(1)}\mid\mathbf{A}_{i}^{(2)};\mathbf{B}^{(2)});\mathbf{B}^{(2),[t]}\right]-p_{\lambda_{N},\tau_{N}}(\mathbf{B}^{(2)}),$ (S.33)
$(\mathbf{B}^{(1),[t+1]},\boldsymbol{\gamma}^{[t+1]}):=$ $\operatorname*{argmax}_{\mathbf{B}^{(1)},\boldsymbol{\gamma}}\sum_{i=1}^{N}\mathbb{E}\left[\log\mathbb{P}(\mathbf{Y}_{i}\mid\mathbf{A}_{i}^{(1)},\mathbf{A}_{i}^{(2)});\mathbf{B}^{(1),[t]},\boldsymbol{\gamma}^{[t]})\right]-p_{\lambda_{N},\tau_{N}}(\mathbf{B}^{(1)}).$ (S.34)
Here, the optimizations for each layer are separated; we update the top-layer proportion parameters $\boldsymbol{p}$} in (S.32), the middle latent layer coefficients $\mathbf{B}^{(2)}$ in (S.33), and the bottom layer coefficients $(\mathbf{B}^{(1)},\boldsymbol{\gamma})$ in (S.34). Additionally, due to the conditional independence assumption in each layer, the maximizations can be further simplified into low-dimensional optimizations over each row of $\mathbf{B}$. Algorithm 5 summarizes the PEM algorithm.
Data: $\mathbf{Y},\mathcal{K}$, tuning parameters $\lambda_{N},\tau_{N}$.
while log-likelihood has not converged do
  In the $[t+1]$th iteration,
 // E-step
  end for
 // M-step
  update $\boldsymbol{\Theta}^{[t+1]}=({\boldsymbol{p}}^{[t+1]},\mathbf{B}^{(1),[t+1]},\mathbf{B}^{(2),[t+1]},\boldsymbol{\gamma}^{[t+1]})$ by solving (S.32)-(S.34).
end while
S.3.2 Detailed Update Formulas for Algorithm 2
Continuing from Section 4, we describe Algorithm 2 under $D=2$ latent layers. The extension to more latent layers is straightforward and we omit the detailed updates for simplicity.
@@
S.3.3 Alternatives for Estimating the Number of Latent Variables
@@
    We postpone the detailed formula of EBIC(k)(k) to the end of this subsection. Here, we consider EBIC instead of more traditional information criteria such as AIC and BIC since it is designed to handle large parameter spaces.
@@
Definition of EBIC.
[…]
S.3.4 Implementation Details
We elaborate on the choices made to implement the EM Algorithms 5 and 2.
@@
Based on the consistency result in Theorem 3, we consider $\lambda_{N}=N^{1/4},\tau_{N}=\max(3N^{-0.3},0.3)$ for the simulation studies. For real data analysis, tuning parameters were selected from the following grid:
@@
In terms of implementing the PEM algorithm to select the number of latent variables, $K^{(1)}$ or $K^{(2)}$, we consider a more generous threshold of $3\times N/500$. This is for the sake of faster implementation, as we only use the resulting likelihood to implement the EBIC and LRT method.
@@
For the simulation studies under deeper models with $D\geq 3$ (see Section 5 in the main paper), we have slightly modified the M-step of the Normal observed-layer (see (S.35)) by considering the hard-thresholding penalty $p_{\tau}(b):=\frac{\tau^{2}}{2}I(|b|\neq 0)$. This leads to a closed-form simplification of (S.35), as spelled-out below. Letting $\mathbf{y}_{j}$ denote the $j$th column of $\mathbf{Y}$, $\mathbf{A}^{[t+1]}_{\text{long}}:=(\mathbf{1}_{N},\mathbf{A}^{(1),[t+1]})$, and $\text{thres}_{\tau}(b):=bI(|b|>\tau)$, we can write the updates for each $\beta_{j},\gamma_{j}$ as
$\widehat{{\boldsymbol{\beta}}}_{j}^{(1),[t+1]}=\text{thres}_{\tau}\Big{(}(\mathbf{A}^{[t+1]\top}_{\text{long}}\mathbf{A}^{[t+1]}_{\text{long}})^{-1}\mathbf{A}^{[t+1]\top}_{\text{long}}\mathbf{y}_{j}\Big{)},\quad\widehat{\gamma}_{j}^{[t+1]}=\sqrt{\frac{\sum_{i}(Y_{i,j}-\eta_{j,\mathbf{A}_{i}^{(1),[t+1]}})^{2}}{N}},$
which speeds up the SAEM algorithm.
For practical implementation for a larger dataset with general observed-layer distributions, we recommend the user to modify the M-step to a faster but approximate optimization procedure. For example, one may choose to do a one-step gradient ascent in each iteration of the M-step.
Addressing latent variable permutation for simulations.
@@
S.4.1 True Parameter Values
In terms of the true parameter values, we consider two sets of values based on the strict and generic identifiability conditions in Theorems 1 and 2, respectively. First, we define $\mathcal{B}_{s}=\{\mathbf{B}_{s}^{(d)}\}_{d\in[D]}$ that satisfy Theorem 1 as follows:
$\mathbf{B}_{s}^{(d)}=\begin{pmatrix}-2\mathbf{1}_{K^{(d)}}&4\mathbf{I}_{K^{(d)}}\\ -4\mathbf{1}_{K^{(d)}}&4\mathbf{I}_{K^{(d)}}\\ -2\mathbf{1}_{K^{(d)}}&\mathbf{B}_{1}^{(d)}\end{pmatrix},\quad\text{where}\quad\beta_{1;j,k}^{(d)}:=\begin{cases}4&\text{if }k=j,\\ 4/3&\text{ else if }|k-j|=K^{(d)}/2,\\ 0&\text{otherwise}.\end{cases}$ (S.36)
We also consider $\mathcal{B}_{g}=\{\mathbf{B}_{g}^{(d)}\}_{d\in[D]}$ that satisfy Theorem 2, defined as:
[…]
and $\mathbf{B}_{1}^{(d)}$ is the matrix defined in Equation S.36. Note that $\mathbf{B}_{g}^{(d)}$ is a less sparse version of $\mathbf{B}_{s}^{(d)}$, where the identity matrices are modified to $\mathbf{B}_{2}^{(d)}$. While $\mathbf{B}_{2}^{(d)}$ and $\mathcal{B}_{g}$ have the same value of 4 on the main diagonals, $\mathcal{B}_{s}$ is sparser. In particular, $\mathcal{B}_{s}$ has two pure children per latent variable whereas $\mathcal{B}_{g}$ has none.
Regarding the remaining parameters, we set the top-layer proportion parameters as $p_{k}=0.5$ for all $k\in[K^{(D)}]$, and set the dispersion parameters for the Normal distribution as $\gamma_{j}=\sigma_{j}^{2}=1$ for all $j\in[J]$.
S.4.2 Additional Details on Ablation Studies
We provide additional details regarding ablation studies. We separately considered data generated from a DDE and a DBN with the same coefficients $\mathcal{B}_{s}$ in (S.36). We implemented the DBN using the Deep Neural Network toolbox in MATLAB (Tanaka and Okutomi, , 2014).
In addition to the analysis in the main text, we compare the RMSE values for the continuous parameters under sparse, identifiable models (panel (d) in Figure 5) versus that under fully-connected, non-identifiable models (Figure S.3 below). Interestingly, the RMSE values returned by the DBN algorithm are similar. This indicates the unreliability of parameter estimation via estimation algorithms for learning DBNs. In contrast, our proposed estimation method for DDEs encourages sparsity as well as adopts a wiser initialization strategy that is close to the true parameters, which leads to better parameter estimation and structure recovery.
Figure S.3: Parameter estimation accuracy for fully-connected DBNs with $J=K^{(1)}=K^{(2)}=18$. The true coefficients are generated iid from $N(0,2)$.
S.4.3 Simulation Results Under the Generic Identifiable Parameters $\mathcal{B}_{g}$ and Computation Time
@@
We present the analogs of Figure 4 under the generically identifiable true parameters $\mathcal{B}_{g}$. While all other simulation settings are identical to that described in the main text, we make the following two changes. First, we do not consider the largest parameter dimension of $(J,K^{(1)},K^{(2)})=(90,30,10)$ in this generically identifiable setting. This is because the layerwise initialization turned out to be less effective in this high-dimensional but less-sparse scenario. Second, for Poisson-based-DDEs, we modify the intercept values to be smaller compared to the values in (S.36) as follows. When $K^{(2)}=2$, we consider a smaller intercept parameter for $\mathbf{B}_{g}^{(1)}$ as follows:
@@
The results, presented in Tables S.6-S.8 (error of estimating $\mathcal{G}$) and S.9-S.11 (error of estimating $\boldsymbol{\Theta}$) in Section S.4.7, show that the PEM estimates exhibit a faster rate of convergence as $N$ increases. In particular, the RMSE values decrease at the parametric rate ${1}/{\sqrt{N}}$, as established in Theorem 3. We believe that the lower estimation accuracy under SAEM is due to multiple approximations within the SAEM algorithm such as approximate sampling and the objective function updates. We recommend using PEM instead of SAEM to estimate DDEs when the latent dimension is small.
@@
S.4.4 Experiments With Varying Numbers of Gibbs Samples in Algorithm 2
@@
[…]
@@
S.4.5 Simulations for Selecting the Number of Latent Variables
@@
Figure S.6: Selection accuracy of $K^{(1)}$ under the two-latent-layer DDE with true parameters $\mathcal{B}_{s}$. The spectral ratio estimator shows near-perfect accuracy for large NN.
@@
We fit the three estimators $400$ times for each scenario, and display the correct selection percentage for the Bernoulli, Poisson, and Normal-based DDEs with true parameters $\mathcal{B}_{s}$$\mathcal{B}_{g}$ in Figure S.6/Figure S.7. Comparing the three estimators, the spectral ratio-based estimator has near-perfect accuracy when $N$ is large enough, say $4000$. This demonstrates the empirical consistency of the spectral ratio estimator. In contrast, the consistency of the EBIC and LRT estimators is not clear for Bernoulli and Poisson responses [^2] To be more precise, as we are using approximate level $\alpha=0.01$ tests, the correct selection percentage of the LRT estimator should converge to 0.99.. Thus, we conclude that for a large enough $N$, it is desirable to select $K^{(1)}$ using the spectral ratio estimator. Among the three response types, the Bernoulli case with its nonlinear link function and limited response values is the most challenging. In contrast, the Normal case with a linear observed layer and continuous responses achieves near-perfect selection accuracy, which is consistent with earlier observations regarding parameter estimation accuracy.
@@
We also apply these three estimators to select $K^{(2)}$, assuming that $K^{(1)}=6$ is correctly estimated or known. The candidate set for $K^{(2)}$ is $\mathfrak{K}:=\{1,2,3\}$. Figure S.8/Figure S.9 displays the correct selection percentage under the true parameters $\mathcal{B}_{s}$$\mathcal{B}_{g}$. Here, we implement the LRT estimator by naively assuming that Wilk’s theorem holds. The overall trends are similar to those observed in the previous figures for selecting $K^{(1)}$. Under the sparse true parameter $\mathcal{B}_{s}$, the spectral ratio estimator performs well across all response types and the LRT estimator has the lowest, but still decent accuracy. Interestingly, the EBIC estimator outperforms the spectral ratio estimator in more challenging scenarios, such as Bernoulli responses with small $N$ or under the Poisson responses with less sparse true parameters $\mathcal{B}_{g}$. Unlike the case for selecting $K^{(1)}$, it is not clear that the spectral ratio estimator outperforms the EBIC estimator for selecting $K^{(2)}$.
@@
S.4.6 Simulations for Deeper Models with $D\geq 3$ Latent Layers
Estimation accuracy of graphical structures
We display the postponed tables from the the main text (Section 5) in Tables S.3 and S.4, which correspond to experiments for deeper models with Normal responses and $D=3,4$ latent layers.
[…]
$\mathbf{G}^{(1)}$ 1.00 1.00 1.00 1.00 1.00 1.00
$\mathbf{G}^{(2)}$ 0.89 0.89 0.90 0.93 0.94 0.98
$\mathbf{G}^{(3)}$ 0.87 0.89 0.88 0.89 0.90 0.91
runtime (s) 3 3 3 5 8 15
Table S.3: Average entrywise-accuracy of estimating graphical matrices and runtime when $D=3$.
[…]
$\mathbf{G}^{(1)}$ 1.00 1.00 1.00 1.00 1.00 1.00
$\mathbf{G}^{(2)}$ 0.83 0.89 0.93 0.95 0.96 0.97
$\mathbf{G}^{(3)}$ 0.72 0.74 0.75 0.77 0.78 0.80
$\mathbf{G}^{(4)}$ 0.64 0.65 0.70 0.70 0.71 0.73
runtime (s) 75 96 74 89 186 312
Table S.4: Average entrywise-accuracy of estimating graphical matrices and runtime when $D=4$.
Selecting the number of latent variables
To assess our spectral-ratio estimator’s performance, we have conducted additional experiments for deeper models with $D=3$. Table S.5 illustrates the satisfactory performance of Algorithm 3 for all layers, where each entry reports the accuracy across 400 replications.We have also evaluated the spectral-ratio estimator for more challenging cases with $D\geq 4$ latent layers, where the accuracy degrades due to accumulation of uncertainty and larger grid size.
[…]
$K^{(1)}$ 100 100 100 100 100 100 14
$K^{(2)}$ 28 44 68 79 87 92 5
$K^{(3)}$ 65 71 82 89 92 95 2
Table S.5: Correct selection percentage for $\mathcal{K}$, under a true model with Normal responses and $D=3,K^{(D)}=2$. The last column reports the typical grid size for each $K^{(d)}$.
S.4.7 Exact Values of Estimation Errors
@@
Tables S.6-S.8 display the average entrywise estimation accuracy for $\mathcal{G}$ under each parametric family. These values correspond to the upper rows of Figures 4 and S.4.
@@
Tables S.9S.11 reports the RMSE values for estimating continuous parameters $\boldsymbol{\Theta}$. These values correspond to the bottom rows of Figures 4 and S.4.
@@
Tables S.12-S.13 reports the accuracy of selecting $\mathcal{K}$ under Bernoulli and Poisson-based DDEs, which correspond to Figure S.6 and Figure S.7. Accuracy values for Normal-based DDEs are omitted, as all methods demonstrated near-perfect accuracy in this setting. Similarly, Tables S.14-S.15 display the results for selecting $K^{(2)}$, which corresponds to Figures S.8 and S.9.
@@
ParFam, value NN Method \ $\widehat{K}^{(1)}$ 4 5 6 7 8
Bernoulli, $\mathcal{B}_{s}$ 500 EBIC 0 0.25 96.25 3.5 0
LRT 0.5 0 98.75 0.75 0
Spectral 1.75 30.0 67.5 0.75 0
1000 EBIC 0 0 94.5 5.5 0
LRT 0 0 96.5 3.5 0
Spectral 0.25 18.25 81.5 0 0
2000 EBIC 0 0 88.25 11.75 0
LRT 0 0 93.5 6.25 0.25
Spectral 0 13.0 87.0 0 0
4000 EBIC 0 0 86.0 12.25 1.75
LRT 0 0 90.25 9.25 0.5
Spectral 0 3.75 96.25 0 0
6000 EBIC 0 0 86.0 13.5 1.5
LRT 0 0 87.75 12.25 0
Spectral 0 0.5 99.5 0 0
8000 EBIC 0 0 87.75 11.0 1.25
LRT 0 0 88.25 0 10.75 1.00
Spectral 0 0.75 99.25 0 0
Bernoulli, $\mathcal{B}_{g}$ 500 EBIC 0 0 84.5 13.75 1.75
LRT 0.75 1.0 86.25 12.0 0
Spectral 8.5 4.25 87.25 0 0
1000 EBIC 0 0 87.25 10.5 2.5
LRT 0 0 80.5 18.5 1.0
Spectral 0.5 0 99.5 0 0
2000 EBIC 0 0 85.25 11.5 3.25
LRT 0 0 82.75 16.25 1.0
Spectral 0 0 100.0 0 0
4000 EBIC 0 0 97.0 3.0 0
LRT 0 0 92.25 7.75 0
Spectral 0 0 100.0 0 0
6000 EBIC 0 0 97.5 2.5 0
LRT 0 0 91.5 8.5 0
Spectral 0 0 100.0 0 0
8000 EBIC 0 0 98.0 2.0 0
LRT 0 0 90.75 9.25 0
Spectral 0 0 100.0 0 0
@@
ParFam, value NN Method \ $\widehat{K}^{(1)}$ 4 5 6 7 8
Poisson, $\mathcal{B}_{s}$ 500 EBIC 0 0 99.5 0.5 0
LRT 0 0 99.5 0.5 0
Spectral 0 0 100.0 0 0
1000 EBIC 0 0 99.0 1.0 0
LRT 0 0 99.0 1.0 0
Spectral 0 0 100.0 0 0
2000 EBIC 0 0 97.5 2.5 0
LRT 0 0 96.25 3.75 0
Spectral 0 0 100.0 0 0
4000 EBIC 0 0 98.25 1.5 0.25
LRT 0 0 97.25 2.75 0
Spectral 0 0 100.0 0 0
6000 EBIC 0 0 99.0 1.0 0
LRT 0 0 97.25 2.75 0
Spectral 0 0 100.0 0 0
8000 EBIC 0 0 98.75 0.75 0.5
LRT 0 0 97.75 1.75 0.5
Spectral 0 0 100.0 0 0
Poisson, $\mathcal{B}_{g}$ 500 EBIC 0 0 92.0 7.75 0.25
LRT 0 1.0 91.0 8.0 0
Spectral 8.5 0 91.5 0 0
1000 EBIC 0 0 90.0 9.5 0.5
LRT 0 0 88.5 10.75 0.75
Spectral 2.75 0 97.25 0 0
2000 EBIC 0 0 90.5 8.5 1.0
LRT 0 0.25 87.25 11.5 1.0
Spectral 0.5 0 99.5 0 0
4000 EBIC 0 0 91.25 7.25 1.5
LRT 0 0 89.0 9.25 1.75
Spectral 0 0 100.0 0 0
6000 EBIC 0 0 95.0 4.5 0.5
LRT 0 0 91.25 7.75 1.0
Spectral 0 0 100.0 0 0
8000 EBIC 0 0 94.0 5.0 1.0
LRT 0 0 90.75 8.25 1.0
Spectral 0 0 100.0 0 0
@@
ParFam, value NN Method \ $\widehat{K}^{(2)}$ 1 2 3
Bernoulli, $\mathcal{B}_{s}$ 500 EBIC 0.5 99.5 0